A 2-liter graduated cylinder initially contains 500ml of water. An irregularly shaped object having a mass of 105gm is submerged into the water. The final volume of the water in the cylinder is 1025ml. Determine the density of the object.
Answer #1

final volume of water in the cylinder =1025ml

Final volume =500ml

rise in volume when the object is added to water= 1025-500= 725ml

105 gm of sample occupies 725ml

density of the sample =105/725ml=0.1448 g/ml